On the flight from Rome to New York , Mona , the flight attendant wheeled her cart which contained 16 cans of Pepsi , 4 cans of coke and 18 cans of ice tea . A passenger asked her to give him a can at random , find the probability that the can is a coke?
Found 3 solutions by Alan3354, jim_thompson5910, rothauserc:
Answer by Alan3354(69443)
Answer by jim_thompson5910(35256)


There are 4 cans of coke. Let's call this value A. So A = 4.

There are 16+4+18 = 38 cans of beverages. Let's call this B, so B = 38

Divide A over B to get A/B = 4/38 = 2/19

Using a calculator, 2/19 = 0.105263 approximately. This converts to 10.5263% after moving the decimal point 2 spots to the right.

-------------------------------------------------------------------------------------------------------------------------------------
-------------------------------------------------------------------------------------------------------------------------------------
The answer as a fraction is 2/19 (which is exact)
The answer as a decimal is 0.105263 (which is approximate; accurate to 6 decimal places)
The answer as a decimal is 10.5263% (which is also approximate)

The answer format you report to the teacher depends on what s/he wants.

Answer by rothauserc(4718)

There are 16 + 4 + 18 = 38 cans in the cart
:
Probability (can selected is coke) = 4/38 = 2/19 is 0.1053 approximately 0.11 or 11%
: